40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 26 16:10:28 2009, Peter King, Configuration, PSL, PSL reference cavity temperature box modifications DSC_0731.jpg
    Reply  Wed Aug 26 16:35:03 2009, Alberto, Configuration, PSL, PSL reference cavity temperature box modifications 
    Reply  Sun May 8 22:59:40 2011, rana, Update, PSL, PSL reference cavity temperature box modifications 
       Reply  Mon May 9 20:30:20 2011, rana, Update, PSL, PSL reference cavity temperature box modifications Untitled.png
          Reply  Tue May 10 20:57:05 2011, rana, Update, PSL, PSL reference cavity temperature box modifications : after 24 hours still OK Untitled.png
Message ID: 4662     Entry time: Sun May 8 22:59:40 2011     In reply to: 1950     Reply to this: 4672
Author: rana 
Type: Update 
Category: PSL 
Subject: PSL reference cavity temperature box modifications 

I looked at the PSL temperature box. It started out as D980400-B-C. Then it was revised by Peter King as per the LHO mods E020247.

There are some more things to do to it to make it useful for us:

  1. R3, 4, 7, 8, 12, & 13 should be changed from 1k to 0 Ohms, I think. I cannot figure out their purpose.
  2. All resistors should be made metal-film. Right now, its kind of a mish-mash.
  3. The active filter U6B has a corner frequency of ~50 Hz. This seems not useful for keeping the 4116 DAC noise out of the temperature. We should lower this to ~30 mHz to take advantage of the stability of the LT1021 which was put in.

** Frank reminds me that we don't use the TIdal or VME external inputs anymore since we moved to the EPICS/Perl PID control. So all we have to do is make sure these inputs are hardware disabled/disconnected.

ELOG V3.1.3-